OpenStudy (anonymous):

An 80.0 kg hiker walks a distance of 400.0 m along a road that slopes 5.0 degrees upward, and then stops. What is the hiker's final gravitational potential energy relative to her original position? 784 J 3.9 × 10^3 J 2.7 × 10^4 J-This is the answer I got, but I am not sure if I am right? 3.1 × 10^5 J

OpenStudy (anonymous):

Sin(5) = h/400 h = 400Sin(5) = 34.8622971 Potential Energy = mgh = mass*gravitational acceleration*height \[PE = (80kg)(9.81m/s^{2})(34.8622971)\] \[= 27,359.93 \frac{ m^2kg }{ s^2 } = 2.7 \times 10^4 J\] So yes, you are right...